# A Physical Model-Guided Framework for Underwater Image Enhancement and Depth Estimation

### v1

@@ Abstract
Due to the selective absorption and scattering of light by diverse aquatic media, underwater images usually suffer from various visual degradations. Existing underwater image enhancement (UIE) approaches that combine underwater physical imaging models with neural networks often fail to accurately estimate imaging model parameters such as depth and veiling light, resulting in poor performance in certain scenarios. To address this issue, we propose a physical model-guided framework for jointly training a Deep Degradation Model (DDM) with any advanced UIE model. DDM includes three well-designed sub-networks to accurately estimate various imaging parameters: a veiling light estimation sub-network, a factors estimation sub-network, and a depth estimation sub-network. Based on the estimated parameters and the underwater physical imaging model, we impose physical constraints on the enhancement process by modeling the relationship between underwater images and desired clean images, i.e., outputs of the UIE model. Moreover, while our framework is compatible with any UIE model, we design a simple yet effective fully convolutional UIE model, termed UIEConv. UIEConv utilizes both global and local features for image enhancement through a dual-branch structure. UIEConv trained within our framework achieves remarkable enhancement results across diverse underwater scenes. Furthermore, as a byproduct of UIE, the trained depth estimation sub-network enables accurate underwater scene depth estimation. Extensive experiments conducted in various real underwater imaging scenarios, including deep-sea environments with artificial light sources, validate the effectiveness of our framework and the UIEConv model.
@@ I Introduction
Underwater images often suffer from color distortion, low contrast, and blurriness caused by light absorption and scattering in water. These issues are exacerbated by varying water conditions such as turbidity and depth which result in uneven lighting and haziness. By improving the visibility and clarity of underwater images, underwater image enhancement (UIE) facilitates the analysis, monitoring, and exploration of underwater scenes.
@@
Early UIE methods simulate the degradation process by an underwater physical imaging model and estimate its parameters to invert clear images [1, 2, 3, 4]. With the introduction of paired datasets containing underwater images and reference images [5, 6, 7], data-driven approaches have gradually gained attention. Numerous UIE models have been proposed to improve performance and efficiency [8, 9, 10, 11, 12]. However, data-driven methods often suffer from poor generalization and interpretability. To combine the advantages of physical model-based methods and the powerful representational capacity of neural networks, some approaches [13, 14] employ neural networks to estimate the parameters of the physical imaging model and then invert the degradation process based on the estimated parameters to obtain enhanced images. We argue that previous methods face two primary issues: (1) inaccurate imaging parameter estimation; and (2) even with accurate parameter estimation, there remains a discrepancy between the physical imaging model and real underwater conditions.
@@
To address the first issue, we design a Deep Degradation Model (DDM), which includes three well-designed sub-networks to estimate scene depth, veiling light, attenuation coefficient, and scattering coefficient. We briefly outline the limitations of previous methods in estimating parameters and our solutions. (1) Scene Depth: To obtain absolute depth, relative depth is typically scaled based on manually set maximum and minimum depths. For example, in HybrUR [14], the minimum and maximum depth in each image are empirically set to 0.1m and 6m, respectively. These predefined depth ranges are unreasonable. Therefore, we propose a depth estimation sub-network to output accurate relative depth maps and a factors estimation sub-network to adaptively output scales and offsets for scaling relative depth values for each image. (2) Veiling Light: Veiling light, considered as the value of backscatter at infinity, is usually represented by a monochrome image where all pixel values are identical across spatial locations [15, 16, 14]. This assumption holds in shallow-water environments where light primarily originates from natural sources but fails in deep-sea scenes with artificial light sources, where underwater images often suffer from uneven lighting [17, 18]. To adapt to various complex environments, we propose a veiling light estimation sub-network that outputs a local background light map where pixel values vary across spatial locations instead of a monochrome image. (3) Attenuation and Scattering Coefficients: When estimating scattering and attenuation coefficients, some methods directly assume they are identical [19]. Even if some methods estimate the transmission map instead of these two coefficients [13, 15, 20], they still rely on the assumption that both coefficients are identical. Recent work [21] has experimentally shown that they are related but not identical. Therefore, our factors estimation sub-network employs two separate heads to distinguish between them, providing a more accurate modeling approach.
@@
To address the second issue, we do not directly use the estimated imaging parameters to invert clean images, as done by previous methods. Instead, we use the physical imaging model to degrade the enhanced image produced by the UIE model. We enforce that the degraded enhanced image closely resembles the original underwater image, thereby imposing physical constraints on the enhancement process of the UIE model. In our training framework, physical constraints manifest as an additional regularization term that assists in the training of the UIE model. In this way, the UIE model with strong fitting capabilities can further be guided and supplemented by domain knowledge of the imaging process. Furthermore, considering that the scene depth of the original and enhanced underwater images should remain consistent, we introduce a depth consistency loss to further assist in the training process. Our framework can improve the performance of many UIE models.
In UIE models, modeling both local and global features is crucial for image restoration. Global features are essential for correcting overall information such as color and lighting, while local features help restore high-frequency details. Due to the limited receptive field of convolutions, some models adopt frequency domain operations [12] or Transformers [7] to extract global features. In contrast, we design a simple yet effective fully convolutional UIE model, termed UIEConv, which includes both a global branch and a local branch. The global branch adopts a U-Net-like architecture, where the encoder progressively downsamples to obtain high-level features, and the decoder upsamples to restore the image. The upsampling and downsampling operations provide a large receptive field, sufficient for modeling long-range dependencies. The local branch, with a smaller receptive field, consists of several convolutions that maintain the image resolution throughout. The final enhanced result is obtained by combining the outputs of the two branches.
@@
  • We test our method in various underwater scenarios, including different water quality environments and deep-sea scenes with limited lighting. There are some examples in Fig. 1. Our approach consistently achieves impressive enhancement and depth estimation results. Extensive experiments validate the effectiveness of our approach.
@@ II-A Underwater Physical Imaging Model
$I$ $=D+B$ (1a)
$=J\cdot e^{-\beta^{D}\cdot d}+B^{\infty}\cdot(1-e^{-\beta^{B}\cdot d}),$ (1b)
where $I\in\mathbb{R}^{3\times H\times W}$ is the observed underwater image, $J\in\mathbb{R}^{3\times H\times W}$ is the restored clear image (i.e., scene radiance), and $B^{\infty}\in\mathbb{R}^{3\times 1\times 1}$ is the veiling light (i.e., global background light). $d$ is the scene depth, representing the distance between the camera and the observed object. Due to the different absorption rates of different wavelengths of light in water, the attenuation coefficient $\beta^{D}$ and the scattering coefficient $\beta^{B}$ are channel-dependent, i.e., $\beta^{D},\beta^{B}\in\mathbb{R}^{3\times 1\times 1}$. As shown on the right side of Fig. 2, red light is absorbed the most in water, resulting in underwater images typically having a bluish-green hue. In some previous literature [2, 3, 15, 19], it was assumed that $\beta^{D}=\beta^{B}=\beta$. Therefore, the underwater physical imaging model can be rewritten as:
$I=J\cdot T+B^{\infty}\cdot(1-T),$ (2)
where $T=e^{-\beta\cdot d}$ is called the transmission map. In this paper, we consider distinguishing between $\beta^{D}$ and $\beta^{B}$ to achieve more accurate modeling. Besides, the veiling light $B^{\infty}$ in Eq. (1b) is assumed to be a monochrome image, i.e., $B^{\infty}\in\mathbb{R}^{3\times 1\times 1}$. This assumption is not reasonable in deep-sea scenarios with artificial light sources where the lighting is typically uneven. In this paper, we address this by assuming that the veiling light may vary at different pixel positions to adapt to scenes with uneven artificial lighting, i.e., $B^{\infty}\in\mathbb{R}^{3\times H\times W}$.
@@ II-B Underwater Image Enhancement
Physical model-free methods directly adjust image pixel values based on various metrics [23]. To improve color cast and contrast degradation in underwater images, some methods perform weighted summation of multiple enhanced versions. For example, Fusion [24] combines the results of white balance and contrast local adaptive histogram equalization. ACCC [25] integrates the complementary advantages of local and global contrast-enhanced versions through multi-scale fusion [26]. MLLE [27] explores adaptive algorithms for color correction and contrast enhancement. Additionally, some methods based on Retinex theory decompose underwater images into reflection and illumination components and enhance them separately [28, 29, 30]. However, these methods often suffer from over-enhancement and color distortions in complex underwater environments.
@@
Physical model-based methods estimate the parameters of the underwater physical imaging model in Eq. (1b) or (2) to invert the degradation process. These parameters can be estimated based on various priors, such as the dark channel prior (DCP) [31, 3], maximum intensity prior [1], image blurriness [4, 32], and minimum information loss [33]. UDCP [2, 34] adapts DCP to underwater environments by using only the blue and green channels to estimate the transmission map. Sea-thru [35] proposes a physically accurate model and restores color based on RGBD images. Berman et al. [36] estimate the attenuation ratios of the blue-red and blue-green color channels by evaluating every possible water type. Zhou et al. [37] estimate depth maps using a channel intensity prior and eliminate backscatter through adaptive dark pixels. However, these methods are sensitive to the assumptions made during parameter estimation, making them less robust in complex and varied underwater environments.
@@
Deep learning-based methods have gained popularity recently due to their remarkable performance. Li et al. [5] introduce the UIEB dataset, which consists of many underwater images and corresponding reference images, for supervised training. This dataset has facilitated the development of various advanced UIE models [11, 38, 12]. Peng et al. [7] construct a lager dataset LSUI and design a Transformer-based UIE model. In addition, UGAN [8] and FUnIEGAN [9] utilize adversarial training to generate clean images. Similar to our work, some methods integrate physical imaging models with neural networks. For instance, some researchers incorporate transmission maps and depth maps as attention or additional inputs to guide neural networks [39, 10]. PUGAN [19] estimates attenuation coefficients and depth maps to derive transmission maps, which then guide the decoding process of recovered images. To circumvent the reliance on paired training data, several self-supervised methods [13, 15, 16, 14] employ neural networks to estimate transmission maps, global background light, and even depth maps, scattering coefficients. These parameters are subsequently used to construct inversion models for image recovery. In comparison to these methods, our approach enables more accurate parameter estimation and better adaptation to various complex underwater environments.
@@
Figure 3: Schematic diagram of our framework. It mainly consists of a deep degradation model (DDM) $\mathcal{H}$ and an underwater image enhancement (UIE) model $\mathcal{F}$. The UIE Model is employed to obtain the enhanced image $\hat{J}$. The DDM is composed of three sub-networks that estimate various parameters of the physical imaging model. Using the outputs from these models, the degraded image $\hat{I}$ can be generated.
@@ III-A Overview of Framework
Our proposed framework is illustrated in Fig. 3. Given an underwater image $I$, the UIE model $\mathcal{F}$ produces an enhanced image $\hat{J}$. The supervised loss $\mathcal{L}_{sup}$ ensures that $\hat{J}$ closely approximates the reference image $J$. Additionally, $I$ is input into a deep degradation model $\mathcal{H}$, which comprises three sub-networks: veiling light estimation sub-network (VLEN), depth estimation network (DEN), and factors estimation sub-network (FEN). VLEN and DEN estimate the veiling light $B^{\infty}$ and relative inverse depth map $z$ of the underwater image $I$, respectively. FEN estimates the attenuation coefficient, scattering coefficient, and the scale factor $z^{Sclae}$ and shift factor $z^{Shift}$ for the relative inverse depth. The two factors $z^{Sclae}$, $z^{Shift}$, and relative inverse depth map $z$ together derive the scene depth map $d$. Finally, all variables are substituted into the underwater physical imaging model in Eq. (1b), resulting in a re-degraded image $\hat{I}$. We enforce consistency between $I$ and $\hat{I}$ to impose a physical constraint loss $\mathcal{L}_{phy}$. Considering that scene depth should be consistent before and after enhancement, we also utilize DEN to estimate the inverse depth map $\hat{z}$ of the enhanced image $\hat{J}$ and introduce an additional depth consistency loss $\mathcal{L}_{depth}$. Ultimately, three losses jointly train the learnable components $\mathcal{H}$ and $\mathcal{F}$ within our framework. We will detail the structures of $\mathcal{H}$ and $\mathcal{F}$ in the following subsections.

### v2

@@ Abstract
Due to the selective absorption and scattering of light by diverse aquatic media, underwater images usually suffer from various visual degradations. Existing underwater image enhancement (UIE) approaches that combine underwater physical imaging models with neural networks often fail to accurately estimate imaging model parameters such as scene depth and veiling light, resulting in poor performance in certain scenarios. To address this issue, we propose a physical model-guided framework for jointly training a Deep Degradation Model (DDM) with any advanced UIE model. DDM includes three well-designed sub-networks to accurately estimate various imaging parameters: a veiling light estimation sub-network, a factors estimation sub-network, and a depth estimation sub-network. Based on the estimated parameters and the underwater physical imaging model, we impose physical constraints on the enhancement process by modeling the relationship between underwater images and desired clean images, i.e., outputs of the UIE model. Moreover, while our framework is compatible with any UIE model, we design a simple yet effective fully convolutional UIE model, termed UIEConv. UIEConv utilizes both global and local features for image enhancement through a dual-branch structure. UIEConv trained within our framework achieves remarkable enhancement results across diverse underwater scenes. Furthermore, as a byproduct of UIE, the trained depth estimation sub-network enables accurate underwater scene depth estimation. Extensive experiments conducted in various real underwater imaging scenarios, including deep-sea environments with artificial light sources, validate the effectiveness of our framework and the UIEConv model.
@@ I Introduction
Underwater images often suffer from color distortion, low contrast, and blurriness caused by light absorption and scattering in water. These issues are further exacerbated by varying water conditions, such as turbidity and depth, which result in uneven lighting and haziness. By improving the visibility and clarity of underwater images, underwater image enhancement (UIE) facilitates the analysis, monitoring, and exploration of underwater scenes.
@@
Early UIE methods simulate the degradation process by an underwater physical imaging model and estimate its parameters to invert clear images [1, 2, 3, 4]. With the introduction of paired datasets containing underwater images and reference images [5, 6, 7], data-driven approaches have gradually gained attention. Numerous UIE models have been proposed to improve performance and efficiency [8, 9, 10, 11, 12]. However, data-driven methods often suffer from poor generalization and interpretability. To combine the advantages of physical model-based methods and the powerful representational capacity of neural networks, some approaches [13, 14] employ neural networks to estimate the parameters of the physical imaging model and then invert the degradation process based on the estimated parameters to obtain enhanced images. We argue that previous methods face two primary issues: (1) inaccurate imaging parameter estimation; and (2) fundamental mismatches between idealized physical imaging models and real-world underwater conditions, making degradation inversion unreliable even with accurate estimation.
TABLE I: Comparison of neural network-based parameter estimation for underwater imaging models. Existing methods either neglect certain parameters (marked N/A) or make simplifying assumptions, while our approach achieves more accurate estimation of all key parameters: scene depth, veiling light, and attenuation/scattering coefficients.
Method Scene Depth Veiling Light Atten./Scatt. Coefficients Coefficients Estimation
USUIR [15] N/A Uniform color image Assumed equal Transmission map
PUGAN [16] Ground truth supervision N/A Assumed equal Single-network
HybrUR [14] CycleGAN loss + rescaled to predefined range Uniform color image Assumed unequal Dual-network
USeReDINet [17] Multi-view synthesis constraints Uniform color image Assumed unequal Dual-network
Proposed Method Physical constraint loss + dynamic rescaling Non-uniform color image Assumed unequal Dual-network
@@
To address the first issue, we design a Deep Degradation Model (DDM), which includes three well-designed sub-networks to estimate scene depth, veiling light, attenuation coefficient, and scattering coefficient. We briefly outline the limitations of previous methods in estimating parameters and our solutions. (1) Scene Depth: To obtain absolute depth, relative depth is typically scaled based on manually set maximum and minimum depths. For example, in HybrUR [14], the minimum and maximum depth in each image are empirically set to 0.1m and 6m, respectively. These predefined depth ranges are unreasonable. Therefore, we propose a depth estimation sub-network to output accurate relative depth maps and a factors estimation sub-network to adaptively output scales and offsets for scaling relative depth values for each image. (2) Veiling Light: Veiling light, considered as the value of backscatter at infinity, is usually represented by a uniform color image where all pixel values are identical across spatial locations [15, 17, 14]. This assumption holds in shallow-water environments where light primarily originates from natural sources but fails in deep-sea scenes with artificial light sources, where underwater images often suffer from uneven lighting [18, 19]. To adapt to various complex environments, we propose a veiling light estimation sub-network that outputs a local background light map where pixel values vary across spatial locations instead of a uniform color image. (3) Attenuation and Scattering Coefficients: When estimating scattering and attenuation coefficients, some methods directly assume they are identical [16]. Even if some methods estimate the transmission map instead of these two coefficients [13, 15], they still rely on the assumption that both coefficients are identical. Recent work [20] has experimentally shown that they are related but not identical. Therefore, our factors estimation sub-network employs two separate heads to distinguish between them, providing a more accurate modeling approach. Table I delineates the distinctions between our method and existing approaches in modeling these imaging parameters.
@@
To address the second issue, we do not directly use the estimated imaging parameters to invert clean images, as done by previous methods. Instead, we use the physical imaging model to degrade the enhanced image produced by the UIE model. We enforce that the degraded enhanced image closely resembles the original underwater image, thereby imposing physical constraints on the enhancement process of the UIE model, as shown in Fig. 3. In our training framework, physical constraints manifest as an additional regularization term that assists in the training of the UIE model. In this way, the UIE model with strong fitting capabilities can further be guided and supplemented by domain knowledge of the imaging process. Furthermore, considering that the scene depth of the original and enhanced underwater images should remain consistent, we introduce a novel depth consistency loss to further assist in the training process.
While our framework is compatible with arbitrary advanced UIE models, we specifically propose a novel UIE model, UIEConv, that offers superior simplicity and effectiveness compared to existing models. Effective image restoration hinges on capturing both local and global features. Global features are critical for correcting overarching aspects such as color and lighting, whereas local features are essential for restoring intricate high-frequency details. Recognizing the inherent limitations of convolutional receptive fields, some models employ frequency domain operations [12] or Transformers [7] to extract global information. Our innovation lies in UIEConv, a simple yet powerful fully convolutional UIE model. It uniquely features distinct global and local branches. The global branch, designed with a U-Net-like architecture, progressively downsamples to capture high-level features and subsequently upsamples to reconstruct the image. This design inherently provides a large receptive field, crucial for modeling long-range dependencies. Complementing this, the local branch, with its more constrained receptive field, utilizes several convolutions to preserve image resolution throughout. The final enhanced result is obtained by combining the outputs of the two branches.
@@
  • We test our method in various underwater scenarios, including different water bodies and deep-sea scenes with limited lighting. As shown in Fig. 1, our approach consistently achieves impressive enhancement and depth estimation results. Extensive experiments validate the effectiveness of our approach.
@@ II-A Underwater Physical Imaging Model
$I_{c}(p)$ $=D_{c}(p)+B_{c}(p)$ (1a)
$=J_{c}(p)\cdot e^{-\beta_{c}^{D}\cdot d(p)}+B^{\infty}_{c}(p)\cdot(1-e^{-\beta^{B}_{c}\cdot d(p)}),$ (1b)
where $c\in\{R,G,B\}$ is the color channel, $p$ is the pixel position, $I\in\mathbb{R}^{3\times H\times W}$ is the observed underwater image, $J\in\mathbb{R}^{3\times H\times W}$ is the restored clear image (i.e., scene radiance), and $B^{\infty}\in\mathbb{R}^{3\times H\times W}$ is the veiling light (i.e., global background light). $I_{c}(p)$ represents the pixel value of the $c$ channel at pixel position $p$ in image $I$. $\beta^{D}_{c},\beta^{B}_{c}\in\mathbb{R}$ are the attenuation and scattering coefficients of the $c$ channel. $d\in\mathbb{R}^{1\times H\times W}$ is the per-pixel scene depth, and $d(p)$ represents the distance from the camera to the corresponding object point $p$ in the scene. As shown on the right side of Fig. 2, red light is absorbed the most in water, resulting in underwater images typically having a bluish-green hue. In some previous literature [2, 3, 15, 16], it was assumed that $\beta^{D}=\beta^{B}=\beta$. Therefore, the underwater physical imaging model can be rewritten as:
$I_{c}(p)=J_{c}(p)\cdot T_{c}(p)+B^{\infty}_{c}(p)\cdot(1-T_{c}(p)),$ (2)
where $T_{c}(p)=e^{-\beta_{c}\cdot d(p)}$ and $T$ is called the transmission map. In this paper, we consider distinguishing between $\beta^{D}$ and $\beta^{B}$ to achieve more accurate modeling. Besides, the veiling light $B^{\infty}$ in Eq. (1b) is assumed to be spatially uniform, implying that it consists of a single and uniform color across the entire image. But this assumption is not reasonable in deep-sea scenarios with artificial light sources where the lighting is typically uneven. In this paper, we address this by assuming that the veiling light may vary at different pixel positions to adapt to scenes with uneven artificial lighting.
@@ II-B Underwater Image Enhancement
Physical model-free methods directly adjust image pixel values based on various metrics [22]. To improve color cast and contrast degradation in underwater images, some methods perform weighted summation of multiple enhanced versions [23]. For example, Fusion [24] combines the results of white balance and contrast local adaptive histogram equalization. ACCC [25] integrates the complementary advantages of local and global contrast-enhanced versions through multi-scale fusion [26]. MLLE [27] explores adaptive algorithms for color correction and contrast enhancement. Additionally, some methods based on Retinex theory decompose underwater images into reflection and illumination components and enhance them separately [28, 29]. However, these methods often suffer from over-enhancement and color distortions in complex underwater environments.
@@
Physical model-based methods estimate the parameters of the underwater physical imaging model in Eq. (1b) or (2) to invert the degradation process. These parameters can be estimated based on various priors, such as the dark channel prior (DCP) [3], maximum intensity prior [1], image blurriness [4, 30], and minimum information loss [31]. UDCP [2, 32] adapts DCP to underwater environments by using only the blue and green channels to estimate the transmission map. Sea-thru [33] proposes a physically accurate model and restores color based on RGBD images. Berman et al. [34] estimate the attenuation ratios of the blue-red and blue-green color channels by evaluating every possible water type. Zhou et al. [35] estimate depth maps using a channel intensity prior and eliminate backscatter through adaptive dark pixels. However, these methods are sensitive to the assumptions made during parameter estimation, making them less robust in complex underwater environments.
@@
Deep learning-based methods have gained popularity recently due to their remarkable performance [36]. Li et al. [5] introduce the UIEB dataset, which consists of many underwater images and corresponding reference images, for supervised training. This dataset has facilitated the development of various advanced UIE models [11, 12, 37]. Peng et al. [7] construct a lager dataset LSUI and design a Transformer-based UIE model. In addition, UGAN [8] and FUnIEGAN [9] utilize adversarial training to generate clean images. Similar to our work, some methods integrate physical imaging models with neural networks. For instance, some researchers incorporate transmission maps and depth maps as attention or additional inputs to guide neural networks [38, 10]. PUGAN [16] estimates attenuation coefficients and depth maps to derive transmission maps, which then guide the decoding process of recovered images. To circumvent the reliance on paired training data, several self-supervised methods [13, 15, 17, 14] employ neural networks to estimate transmission maps, global background light, and even depth maps, scattering coefficients. These parameters are subsequently used to construct inversion models for image recovery. In comparison to these methods, our approach enables more accurate parameter estimation and better adaptation to various complex underwater environments.
Figure 3: Schematic diagram of our physical model-guided framework. It primarily consists of a deep degradation model (DDM) $\mathcal{H}$ and an underwater image enhancement (UIE) model $\mathcal{F}$. The UIE Model is used to obtain the enhanced image $\hat{J}$. The DDM is composed of three sub-networks that estimate various parameters of the imaging model. Using the outputs from $\mathcal{H}$ and $\mathcal{F}$, the re-degraded image $\hat{I}$ can be generated by the underwater physical imaging model. For simplicity, we omit the channel $c$ and pixel position $p$ in the equation.
@@ III-A Overview of Framework
Our proposed framework is illustrated in Fig. 3. Given an underwater image $I$, the UIE model $\mathcal{F}$ produces an enhanced image $\hat{J}$. The supervised loss $\mathcal{L}_{sup}$ ensures that $\hat{J}$ closely approximates the reference image $J$. Additionally, $I$ is input into a Deep Degradation Model (DDM) $\mathcal{H}$, which comprises three sub-networks: veiling light estimation sub-network (VLEN), depth estimation network (DEN), and factors estimation sub-network (FEN). VLEN and DEN estimate the veiling light $B^{\infty}$ and relative inverse depth map $z$ of the underwater image $I$, respectively. FEN estimates the attenuation coefficient, scattering coefficient, and the scale factor $z^{Sclae}$ and shift factor $z^{Shift}$ for the relative inverse depth. The two factors $z^{Sclae}$, $z^{Shift}$, and relative inverse depth map $z$ together derive the scene depth map $d$. Finally, all variables are substituted into the underwater physical imaging model in Eq. (1b), resulting in a re-degraded image $\hat{I}$. We enforce consistency between $I$ and $\hat{I}$ to impose a physical constraint loss $\mathcal{L}_{phy}$. It is noteworthy that for the parameters of the underwater physical imaging model output by DDM, such as scene depth $d$ and scattering coefficients $\beta^{B}$, we do not have their ground-truth values for supervised training. Therefore, we ingeniously utilize the physical constraint loss $\mathcal{L}_{phy}$, derived from the relationship between the raw and re-degraded images, to guide the DDM toward accurate estimations. In turn, these estimated parameters provide physical guidance for the training of the UIE model. Besides, considering that scene depth should be consistent before and after enhancement, we also utilize DEN to estimate the inverse depth map $\hat{z}$ of the enhanced image $\hat{J}$ and introduce an additional depth consistency loss $\mathcal{L}_{depth}$. Ultimately, three losses jointly train the learnable components $\mathcal{H}$ and $\mathcal{F}$ within our framework. We will detail the structures of $\mathcal{H}$ and $\mathcal{F}$ in the following subsections.